# Determination of Magnetic Symmetries by Convergent Beam Electron Diffraction

### v2

@@ VII Summary
In this work, we have introduced magnetic CBED for the unambiguous determination of magnetic-point-group symmetries in magnetic crystalline materials. We have classified all magnetic point-group symmetries of the diffraction patterns in terms of 125 magnetic CBED groups. The subsequent mapping of all magnetic point groups to corresponding magnetic CBED groups allows the determination of the latter from a finite series of CBED experiments on slabs of different crystal orientation. The signal level (electron dose) required for a successful experimental realization of the proposed method is achievable in modern TEMs. First experimental realizations of this measurement principle have been demonstrated for ferromagnets previously. Simplicity and cost-effectiveness make magnetic CBED an intriguing alternative to neutron diffraction for magnetic-symmetry determination. The high spatial sensitivity of magnetic CBED, allowing to probe very small volumes down to about $10\,\mathrm{nm}\times 10\,\mathrm{nm}\times 100\,\mathrm{nm}$, should facilitate the probing of magnetic symmetries for the large class of magnetic micro- and nanoparticles, polycrystalline magnetic matter, magnetic thin films, and magnetic van der Waals materials, that cannot be measured with neutron diffraction. Moreover, the high spatial resolution down to about $10\,\mathrm{nm}$ in principle facilitates mapping of symmetries across phase and domain boundaries, or close to crystal defects, by recording magnetic CBED patterns at different beam positions. We therefore foresee manifold applications for magnetic CBED as a symmetry probe for magnetic crystalline matter.
@@ Appendix A Detailed derivation of the paraxial scattering equation
Because of to high energy $E$ of highly accelerated electrons in a Transmission Electron Microscope the electric potential squared term can be neglected, which is referred to as the high-energy approximation [62]. After a couple of rearrangements and dividing by $E=\gamma m_{e}c^{2}$, where $\gamma=1/\sqrt{1-v^{2}/c^{2}}$ is the Lorentz factor, this high-energy limit of the Klein-Gordon equation takes a form that is mathematically equivalent to the stationary Schrödinger equation [28]
@@ Appendix B Details of the construction of magnetic diffraction groups
Let $G$ denote any of the structural point groups $C_{n}$ and $D_{n}$ with $n=1,2,3,4,6$, where we use the standard Schoenflies notation for two-dimensional point groups (rosette groups). Adding $\mathcal{T}$ as a new generator leads to the $10$ groups $G+G\mathcal{T}$, which we now call “$\mathcal{T}$ gray” to distinguish them from the following class. Instead adding $\mathcal{Z}$ leads to the $10$ groups $G+G\mathcal{Z}$, which we call “$\mathcal{Z}$ gray.” Finally, adding $\Omega$ as a new generator leads to the $10$ groups $G+G\Omega$. These groups only contain unitary elements and are thus monochromatic. We can also add two of $\mathcal{T}$, $\mathcal{Z}$, or $\Omega$ as new generators. Since the product of any two of these is the third, this only produces $10$ additional groups $G+G\mathcal{T}+G\mathcal{Z}+G\Omega$. These groups are both $\mathcal{T}$ gray and $\mathcal{Z}$ gray. The groups constructed so far are listed in Tables 2 and 3.
Table 2: Structural monochromatic (first and second column), $\mathcal{T}$ gray (third column), and $\mathcal{Z}$ gray (fourth column) CBED groups. The Schoenflies notation is given for the monochromatic groups. The Hermann–Mauguin notation [63] is presented for all groups, where $1^{\prime}$ refers to the time-reversal operation $\mathcal{T}$ and $1_{R}$ denotes the z-reversal operation $\mathcal{Z}$.
@@
$C_{1}$ $1$ $11^{\prime}$ $11_{R}$
$C_{2}$ $2$ $21^{\prime}$ $21_{R}$
$C_{3}$ $3$ $31^{\prime}$ $31_{R}$
$C_{4}$ $4$ $41^{\prime}$ $41_{R}$
$C_{6}$ $6$ $61^{\prime}$ $61_{R}$
@@
Table 3: Monochromatic CBED groups containing the product $\Omega=\mathcal{T}\mathcal{Z}$ of time and z reversal (first column) and CBED groups that are both $\mathcal{T}$ gray and $\mathcal{Z}$ gray (second column). The Hermann–Mauguin notation [63] is presented for all groups. $1^{\prime}_{R}$ denotes the product operation $\Omega$. Note that $1_{R}^{\prime}$ is not given explicitly for the $\mathcal{T}$ and $\mathcal{Z}$ gray groups since it is generated by $1^{\prime}$ and $1_{R}$.
@@
Table 4: Dichromatic CBED groups based on time reversal $\mathcal{T}$ (first and second column). The Schoenflies notation $G(H)$ for the group $H+(G\setminus H)\mathcal{T}$ and the Hermann–Mauguin notation [63] are given. The third and fourth column show the $\mathcal{Z}$ dichromatic CBED groups and the “pseudo-dichromatic” (but actually monochromatic) CBED groups based on $\Omega$. In the Hermann–Mauguin notation, a prime means that the corresponding generator is multiplied by the time-reversal operation $\mathcal{T}$, a subscript “R” denotes multiplication by $\mathcal{Z}$, and a prime combined with a subscript “R” denotes multiplication by $\Omega$.
@@ Appendix C Mappings of magnetic point groups to magnetic CBED groups
In this Appendix, we list the mapping from the 122 three-dimensional magnetic point groups to the 125 magnetic CBED groups for all distinct beam directions. The tables are organized according to crystal classes: The mappings for triclinic point groups are shown in Table 7, those for monoclinic groups in Table 8, those for orthorhomic groups in Table 9, those for tetragonal groups in Tables 10 and 11, those for trigonal groups in Tables 12 and 13, those for hexagonal groups in Tables 14 and 15, and those for cubic groups in Tables 16 and 17. In all tables, the mapping of magnetic point groups (mPG) to magnetic CBED groups (magnetic diffraction groups, mDG) is shown for all slab orientations. The underlying structural point group is given in Hermann-Mauguin and, at the first occurrence, also in Schoenflies notation. The slab normal is parallel to the beam direction, which is denoted by $[uvw]$ in crystallographic coordinates. Here, variables $u$, $v$, $w$ stand for generic values, which are distinct from the special values given in the tables to denote high-symmetry directions. All tables also show the corresponding structural point group (PG). We note that there are occasional permutations of the symmetry symbols in the point group notations, e.g., $mm2$ instead of $2mm$, corresponding to different choices of unit cell orientations. By using them we follow established conventions in the literature [12, 13]. These permutations do not affect the symmetry considerations.
@@
$2$ ($C_{2}$) $2$ $2$ $m_{R}$ $1$
$2$ $21^{\prime}$ $21^{\prime}$ $m_{R}1^{\prime}$ $11^{\prime}$
$2$ $2^{\prime}$ $2^{\prime}$ $m^{\prime}_{R}$ $1$
$m$ ($C_{s}$) $m$ $1_{R}$ $m$ $1$
@@
$m$ $m^{\prime}$ $1^{\prime}_{R}$ $m^{\prime}$ $1$
@@
$222$ $22^{\prime}2^{\prime}$ 2 $2^{\prime}m_{R}m^{\prime}_{R}$ $2m^{\prime}_{R}m^{\prime}_{R}$ $2^{\prime}m_{R}m^{\prime}_{R}$ $m^{\prime}_{R}$ $m_{R}$ $m^{\prime}_{R}$ $1$
$mm2$ ($C_{2v}$) $mm2$ $m1_{R}$ $2mm$ $m1_{R}$ $m$ $m_{R}$ $m$ $1$
$mm2$ $mm21^{\prime}$ $m1_{R}1^{\prime}$ $2mm1^{\prime}$ $m1_{R}1^{\prime}$ $m1^{\prime}$ $m_{R}1^{\prime}$ $m1^{\prime}$ $11^{\prime}$
$mm2$ $m^{\prime}m^{\prime}2$ $m^{\prime}1^{\prime}_{R}$ $2m^{\prime}m^{\prime}$ $m^{\prime}1^{\prime}_{R}$ $m^{\prime}$ $m_{R}$ $m^{\prime}$ $1$
$mm2$ $m^{\prime}m2^{\prime}$ $m^{\prime}1_{R}$ $2^{\prime}mm^{\prime}$ $m1^{\prime}_{R}$ $m$ $m^{\prime}_{R}$ $m^{\prime}$ $1$
$mmm$ ($D_{2h}$) $mmm$ $2mm1_{R}$ $2mm1_{R}$ $2mm1_{R}$ $2_{R}mm_{R}$ $2_{R}mm_{R}$ $2_{R}mm_{R}$ $2_{R}$
$mmm$ $mmm1^{\prime}$ $2mm1_{R}1^{\prime}$ $2mm1_{R}1^{\prime}$ $2mm1_{R}1^{\prime}$ $2_{R}mm_{R}1^{\prime}$ $2_{R}mm_{R}1^{\prime}$ $2_{R}mm_{R}1^{\prime}$ $2_{R}1^{\prime}$
$mmm$ $m^{\prime}m^{\prime}m$ $2^{\prime}mm^{\prime}1^{\prime}_{R}$ $2m^{\prime}m^{\prime}1_{R}$ $2^{\prime}mm^{\prime}1^{\prime}_{R}$ $2_{R}m^{\prime}m^{\prime}_{R}$ $2_{R}mm_{R}$ $2_{R}m^{\prime}m^{\prime}_{R}$ $2_{R}$
$mmm$ $m^{\prime}m^{\prime}m^{\prime}$ $2m^{\prime}m^{\prime}1^{\prime}_{R}$ $2m^{\prime}m^{\prime}1^{\prime}_{R}$ $2m^{\prime}m^{\prime}1^{\prime}_{R}$ $2^{\prime}_{R}m^{\prime}m_{R}$ $2^{\prime}_{R}m^{\prime}m_{R}$ $2^{\prime}_{R}m^{\prime}m_{R}$ $2^{\prime}_{R}$
$mmm$ $mmm^{\prime}$ $2^{\prime}mm^{\prime}1_{R}$ $2mm1^{\prime}_{R}$ $2^{\prime}mm^{\prime}1_{R}$ $2^{\prime}_{R}mm^{\prime}_{R}$ $2^{\prime}_{R}m^{\prime}m_{R}$ $2^{\prime}_{R}mm^{\prime}_{R}$ $2^{\prime}_{R}$
@@
$4$ ($C_{4}$) $4$ $4$ $m_{R}$ $1$
$4$ $41^{\prime}$ $41^{\prime}$ $m_{R}1^{\prime}$ $11^{\prime}$
$4$ $4^{\prime}$ $4^{\prime}$ $m_{R}$ $1$
$\bar{4}$ ($S_{4}$) $\bar{4}$ $4_{R}$ $m_{R}$ $1$
@@
$422$ $4^{\prime}22^{\prime}$ $4^{\prime}m_{R}m^{\prime}_{R}$ $2m_{R}m_{R}$ $2^{\prime}m_{R}m^{\prime}_{R}$ $m_{R}$ $m_{R}$ $m^{\prime}_{R}$ $1$
$422$ $42^{\prime}2^{\prime}$ $4m^{\prime}_{R}m^{\prime}_{R}$ $2^{\prime}m_{R}m^{\prime}_{R}$ $2^{\prime}m_{R}m^{\prime}_{R}$ $m_{R}^{\prime}$ $m_{R}$ $m^{\prime}_{R}$ $1$
$4mm$ ($C_{4v}$) $4mm$ $4mm$ $m1_{R}$ $m1_{R}$ $m$ $m_{R}$ $m$ $1$
$4mm$ $4mm1^{\prime}$ $4mm1^{\prime}$ $m1_{R}1^{\prime}$ $m1_{R}1^{\prime}$ $m1^{\prime}$ $m_{R}1^{\prime}$ $m1^{\prime}$ $11^{\prime}$
$4mm$ $4^{\prime}mm^{\prime}$ $4^{\prime}mm^{\prime}$ $m1_{R}$ $m^{\prime}1^{\prime}_{R}$ $m$ $m_{R}$ $m^{\prime}$ $1$
$4mm$ $4m^{\prime}m^{\prime}$ $4m^{\prime}m^{\prime}$ $m^{\prime}1^{\prime}_{R}$ $m^{\prime}1^{\prime}_{R}$ $m^{\prime}$ $m_{R}$ $m^{\prime}$ $1$
$\bar{4}2m$ ($D_{2d}$) $\bar{4}2m$ $4_{R}mm_{R}$ $2m_{R}m_{R}$ $m1_{R}$ $m_{R}$ $m_{R}$ $m$ $1$
@@
$\bar{4}2m$ $\bar{4}^{\prime}2m^{\prime}$ $4^{\prime}_{R}m^{\prime}m_{R}$ $2m_{R}m_{R}$ $m^{\prime}1^{\prime}_{R}$ $m_{R}$ $m_{R}$ $m^{\prime}$ $1$
$\bar{4}2m$ $\bar{4}^{\prime}2^{\prime}m$ $4^{\prime}_{R}mm^{\prime}_{R}$ $2^{\prime}m_{R}m^{\prime}_{R}$ $m1_{R}$ $m^{\prime}_{R}$ $m_{R}$ $m$ $1$
$\bar{4}2m$ $\bar{4}2^{\prime}m^{\prime}$ $4_{R}m^{\prime}m^{\prime}_{R}$ $2^{\prime}m_{R}m^{\prime}_{R}$ $m^{\prime}1^{\prime}_{R}$ $m^{\prime}_{R}$ $m_{R}$ $m^{\prime}$ $1$
@@
$3$ ($C_{3}$) $3$ $3$ $1$
$3$ $31^{\prime}$ $31^{\prime}$ $11^{\prime}$

### v3

@@ VII Summary
In this work, we have introduced magnetic CBED for the unambiguous determination of magnetic-point-group symmetries in magnetic crystalline materials. We have classified all magnetic point-group symmetries of the diffraction patterns in terms of 125 magnetic CBED groups. The subsequent mapping of all magnetic point groups to corresponding magnetic CBED groups allows the determination of the latter from a finite series of CBED experiments on slabs of different crystal orientation. A successful experimental realization of the magnetic CBED requires overcoming several experimental challenges, including artificial symmetry breaking by beam damage and adsorbates. Our calculation indicate, however, that the signal level (electron dose) required for a successful experimental realization of the proposed method for elucidating the symmetries in antiferromagnets is achievable in modern TEMs. First experimental realizations of this measurement principle have been demonstrated for ferromagnets previously.
Theoretically, simplicity and cost-effectiveness should make magnetic CBED an intriguing alternative to neutron diffraction for magnetic-symmetry determination. The high spatial sensitivity of magnetic CBED, allowing to probe very small volumes down to about $10\,\mathrm{nm}\times 10\,\mathrm{nm}\times 100\,\mathrm{nm}$, should facilitate the probing of magnetic symmetries for magnetic micro- and nanoparticles, polycrystalline magnetic matter, magnetic thin films, and magnetic van der Waals materials, that cannot be measured with neutron diffraction. Moreover, the high spatial resolution down to about $10\,\mathrm{nm}$ in principle facilitates mapping of symmetries across phase and domain boundaries, or close to crystal defects, by recording magnetic CBED patterns at different beam positions. As a note of caution, we again emphasize that all these considerations hinge particularly on the beam stability of the sample that is required to achieve sufficient signal to noise ratio for probing the magnetic symmetry.
@@ Appendix A Detailed derivation of the paraxial scattering equation
Because of to high energy $E$ of highly accelerated electrons in a Transmission Electron Microscope the electric potential squared term can be neglected, which is referred to as the high-energy approximation [Fujiwara1961]. After a couple of rearrangements and dividing by $E=\gamma m_{e}c^{2}$, where $\gamma=1/\sqrt{1-v^{2}/c^{2}}$ is the Lorentz factor, this high-energy limit of the Klein-Gordon equation takes a form that is mathematically equivalent to the stationary Schrödinger equation [Rother2009]
@@ Appendix B Details of the construction of magnetic diffraction groups
Let $G$ denote any of the structural point groups $C_{n}$ and $D_{n}$ with $n=1,2,3,4,6$, where we use the standard Schoenflies notation for two-dimensional point groups (rosette groups). Adding $\mathcal{T}$ as a new generator leads to the $10$ groups $G+G\mathcal{T}$, which we now call $\mathcal{T}$ gray” to distinguish them from the following class. Instead adding $\mathcal{Z}$ leads to the $10$ groups $G+G\mathcal{Z}$, which we call $\mathcal{Z}$ gray.” Finally, adding $\Omega$ as a new generator leads to the $10$ groups $G+G\Omega$. These groups only contain unitary elements and are thus monochromatic. We can also add two of $\mathcal{T}$, $\mathcal{Z}$, or $\Omega$ as new generators. Since the product of any two of these is the third, this only produces $10$ additional groups $G+G\mathcal{T}+G\mathcal{Z}+G\Omega$. These groups are both $\mathcal{T}$ gray and $\mathcal{Z}$ gray. The groups constructed so far are listed in Tables 2 and 3.
Table 2: Structural monochromatic (first and second column), $\mathcal{T}$ gray (third column), and $\mathcal{Z}$ gray (fourth column) CBED groups. The Schoenflies notation is given for the monochromatic groups. The Hermann–Mauguin notation [HM_notation] is presented for all groups, where $1^{\prime}$ refers to the time-reversal operation $\mathcal{T}$ and $1_{R}$ denotes the z-reversal operation $\mathcal{Z}$.
@@
$C_{1}$ 11 $1^{\prime} 11$$1_{R}$
$C_{2}$ 22 $21^{\prime}$ $21_{R}$
$C_{3}$ 33 $31^{\prime}$ $31_{R}$
$C_{4}$ 44 $41^{\prime}$ $41_{R}$
$C_{6}$ 66 $61^{\prime}$ $61_{R}$
@@
Table 3: Monochromatic CBED groups containing the product $\Omega=\mathcal{T}\mathcal{Z}$ of time and z reversal (first column) and CBED groups that are both $\mathcal{T}$ gray and $\mathcal{Z}$ gray (second column). The Hermann–Mauguin notation [HM_notation] is presented for all groups. $1^{\prime}_{R}$ denotes the product operation $\Omega$. Note that $1_{R}^{\prime}$ is not given explicitly for the $\mathcal{T}$ and $\mathcal{Z}$ gray groups since it is generated by $1^{\prime}$ and $1_{R}$.
@@
Table 4: Dichromatic CBED groups based on time reversal $\mathcal{T}$ (first and second column). The Schoenflies notation $G(H)$ for the group $H+(G\setminus H)\mathcal{T}$ and the Hermann–Mauguin notation [HM_notation] are given. The third and fourth column show the $\mathcal{Z}$ dichromatic CBED groups and the “pseudo-dichromatic” (but actually monochromatic) CBED groups based on $\Omega$. In the Hermann–Mauguin notation, a prime means that the corresponding generator is multiplied by the time-reversal operation $\mathcal{T}$, a subscript “R” denotes multiplication by $\mathcal{Z}$, and a prime combined with a subscript “R” denotes multiplication by $\Omega$.
@@ Appendix C Mappings of magnetic point groups to magnetic CBED groups
In this Appendix, we list the mapping from the 122 three-dimensional magnetic point groups to the 125 magnetic CBED groups for all distinct beam directions. The tables are organized according to crystal classes: The mappings for triclinic point groups are shown in Table 7, those for monoclinic groups in Table 8, those for orthorhomic groups in Table 9, those for tetragonal groups in Tables 10 and 11, those for trigonal groups in Tables 12 and 13, those for hexagonal groups in Tables 14 and 15, and those for cubic groups in Tables 16 and 17. In all tables, the mapping of magnetic point groups (mPG) to magnetic CBED groups (magnetic diffraction groups, mDG) is shown for all slab orientations. The underlying structural point group is given in Hermann-Mauguin and, at the first occurrence, also in Schoenflies notation. The slab normal is parallel to the beam direction, which is denoted by $[uvw]$ in crystallographic coordinates. Here, variables $u$, $v$, $w$ stand for generic values, which are distinct from the special values given in the tables to denote high-symmetry directions. All tables also show the corresponding structural point group (PG). We note that there are occasional permutations of the symmetry symbols in the point group notations, e.g., $mm2$ instead of $2mm$, corresponding to different choices of unit cell orientations. By using them we follow established conventions in the literature [Buxton1976, Tanaka2010]. These permutations do not affect the symmetry considerations.
@@ 11 $^{\prime} 1111^{\prime}$
22 ($C_{2}$) 22 22 $m_{R}$ 11
22 $^{\prime} 2121^{\prime}$ $m_{R}1^{\prime}$ $11^{\prime}$
22 $^{\prime} 22^{\prime}$ $m^{\prime}_{R}$ 11
$m$ ($C_{s}$) $m$ $1_{R}$ $m$ 11
@@ 22 $^{\prime} 22^{\prime}$ $m^{\prime}_{R}$ 11
$m$ $m^{\prime}$ $1^{\prime}_{R}$ $m^{\prime}$ 11
@@
$222$ $2^{\prime}2^{\prime}2$ $2^{\prime}m_{R}m^{\prime}_{R}$ $2m^{\prime}_{R}m^{\prime}_{R}$ $2^{\prime}m_{R}m^{\prime}_{R}$ $m^{\prime}_{R}$ $m_{R}$ $m^{\prime}_{R}$ 11
$mm2$ ($C_{2v}$) $mm2$ $m1_{R}$ $2mm$ $1_{R} mm$ $_{R} mm$ 11
$mm2$ $mm21^{\prime}$ $m1_{R}1^{\prime}$ $2mm1^{\prime}$ $_{R}1^{\prime} m1m1^{\prime}$ $m_{R}1^{\prime}$ $m1^{\prime}$ $11^{\prime}$
$mm2$ $m^{\prime}m^{\prime}2$ $m^{\prime}1^{\prime}_{R}$ $2m^{\prime}m^{\prime}$ $^{\prime}1^{\prime}_{R} mm^{\prime}$ $_{R} mm^{\prime}$ 11
$mm2$ $m^{\prime}m2^{\prime}$ $m^{\prime}1_{R}$ $2^{\prime}mm^{\prime}$ $1^{\prime}_{R} mm$ $^{\prime}_{R} mm^{\prime}$ 11
mmmmmm ($D_{2h}$) mmmmmm $2mm1_{R}$ $2mm1_{R}$ $2mm1_{R}$ $2_{R}mm_{R}$ $2_{R}mm_{R}$ $2_{R}mm_{R}$ $2_{R}$
mmmmmm $mmm1^{\prime}$ $2mm1_{R}1^{\prime}$ $2mm1_{R}1^{\prime}$ $2mm1_{R}1^{\prime}$ $2_{R}mm_{R}1^{\prime}$ $2_{R}mm_{R}1^{\prime}$ $2_{R}mm_{R}1^{\prime}$ $2_{R}1^{\prime}$
mmmmmm $m^{\prime}m^{\prime}m$ $2^{\prime}mm^{\prime}1^{\prime}_{R}$ $2m^{\prime}m^{\prime}1_{R}$ $2^{\prime}mm^{\prime}1^{\prime}_{R}$ $2_{R}m^{\prime}m^{\prime}_{R}$ $2_{R}mm_{R}$ $2_{R}m^{\prime}m^{\prime}_{R}$ $2_{R}$
mmmmmm $m^{\prime}m^{\prime}m^{\prime}$ $2m^{\prime}m^{\prime}1^{\prime}_{R}$ $2m^{\prime}m^{\prime}1^{\prime}_{R}$ $2m^{\prime}m^{\prime}1^{\prime}_{R}$ $2^{\prime}_{R}m^{\prime}m_{R}$ $2^{\prime}_{R}m^{\prime}m_{R}$ $2^{\prime}_{R}m^{\prime}m_{R}$ $2^{\prime}_{R}$
mmmmmm $mmm^{\prime}$ $2^{\prime}mm^{\prime}1_{R}$ $2mm1^{\prime}_{R}$ $2^{\prime}mm^{\prime}1_{R}$ $2^{\prime}_{R}mm^{\prime}_{R}$ $2^{\prime}_{R}m^{\prime}m_{R}$ $2^{\prime}_{R}mm^{\prime}_{R}$ $2^{\prime}_{R}$
@@
44 ($C_{4}$) 44 44 $m_{R}$ 11
44 $^{\prime} 4141^{\prime}$ $m_{R}1^{\prime}$ $11^{\prime}$
44 $^{\prime} 44^{\prime}$ $m_{R}$ 11
$\bar{4}$ ($S_{4}$) $\bar{4}$ $4_{R}$ $m_{R}$ 11
@@ 44 $^{\prime} 44^{\prime}$ $m_{R}$ 11
$422$ $4^{\prime}22^{\prime}$ $4^{\prime}m_{R}m^{\prime}_{R}$ $2m_{R}m_{R}$ $2^{\prime}m_{R}m^{\prime}_{R}$ $m_{R}$ $m_{R}$ $m^{\prime}_{R}$ 11
$422$ $42^{\prime}2^{\prime}$ $4m^{\prime}_{R}m^{\prime}_{R}$ $2^{\prime}m_{R}m^{\prime}_{R}$ $2^{\prime}m_{R}m^{\prime}_{R}$ $m_{R}^{\prime}$ $m_{R}$ $m^{\prime}_{R}$ 11
$4mm$ ($C_{4v}$) $4mm$ $4mm$ $m1_{R}$ $1_{R} mm$ $_{R} mm$ 11
$4mm$ $^{\prime} 4mm14mm1^{\prime}$ $m1_{R}1^{\prime}$ $_{R}1^{\prime} m1m1^{\prime}$ $m_{R}1^{\prime}$ $m1^{\prime}$ $11^{\prime}$
$4mm$ $4^{\prime}mm^{\prime}$ $4^{\prime}mm^{\prime}$ $m1_{R}$ $^{\prime}1^{\prime}_{R} mm$ $_{R} mm^{\prime}$ 11
$4mm$ $^{\prime}m^{\prime} 4mm4$$^{\prime}m^{\prime}$ $m^{\prime}1^{\prime}_{R}$ $^{\prime}1^{\prime}_{R} mm^{\prime}$ $_{R} mm^{\prime}$ 11
$\bar{4}2m$ ($D_{2d}$) $\bar{4}2m$ $4_{R}mm_{R}$ $2m_{R}m_{R}$ $m1_{R}$ $m_{R}$ $_{R} mm$ 11
@@
$\bar{4}2m$ $\bar{4}^{\prime}2m^{\prime}$ $4^{\prime}_{R}m^{\prime}m_{R}$ $2m_{R}m_{R}$ $m^{\prime}1^{\prime}_{R}$ $m_{R}$ $_{R} mm^{\prime}$ 11
$\bar{4}2m$ $\bar{4}^{\prime}2^{\prime}m$ $4^{\prime}_{R}mm^{\prime}_{R}$ $2^{\prime}m_{R}m^{\prime}_{R}$ $m1_{R}$ $m^{\prime}_{R}$ $_{R} mm$ 11
$\bar{4}2m$ $\bar{4}2^{\prime}m^{\prime}$ $4_{R}m^{\prime}m^{\prime}_{R}$ $2^{\prime}m_{R}m^{\prime}_{R}$ $m^{\prime}1^{\prime}_{R}$ $m^{\prime}_{R}$ $_{R} mm^{\prime}$ 11
@@
33 ($C_{3}$) 33 33 11
33 $^{\prime} 3131^{\prime}$ $11^{\prime}$